OpenStudy (anonymous):

The volume of a cube is found by using the formula , where l is the side length. If the side length is , what is the volume of the cube?

OpenStudy (anonymous):

f the side lenghth is an variable x then the volume of the cube will be x*x*x=x^3

OpenStudy (anonymous):

example a sidelenghth of a cube is 2 so the volume of the cube will be 2*2*2=2^3=8

OpenStudy (anonymous):

THATS WRONG THAT NOT PART OF MY ANSWER CHOICES